A truck gradually starts off from rest with a uniform acceleration of 2 m/s2 . It reaches a velocity of 16 m/s. Using the equation v2 = u2 + 2as, calculate the distance it travels while it is accelerating.


1
Expert's answer
2021-03-24T19:50:15-0400
"v^2=u^2+2as,""s=\dfrac{v^2-u^2}{2a},""s=\dfrac{(16\ \dfrac{m}{s})^2-0}{2\cdot2\ \dfrac{m}{s^2}}=64\ m."